OpenStudy (anonymous):

Help please? I'll give you a medal .. The formula for degrees Celsius is: C = 5/9 (F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is.

OpenStudy (ivettef365):

to solve for F, you need to isolate the F as follows: C = 5/9(F-32) (9/5) C = 5/9 (F-32) (9/5) multiply both side by 9 then you have 9/5C = F - 32 (THEN ADD BOTH SIDE BY 32) 9/5C + 32 = F so you have that F = 9/5C+32

OpenStudy (ivettef365):

I meant (multiply both sides by 9/5)

OpenStudy (ivettef365):

the second just substitute 6 for C

OpenStudy (anonymous):

thank you ;)